# Parasitic RF-SQUIDs in superconducting qubits due to wirebonds

### v1

@@ Introduction
The qubit transition frequency $f_{\mathrm{Q}}$ depends on the flux $\Phi_{\mathrm{Q}}$ threading its DC-SQUID and can be approximated by[12]
@@
Figure 3: a Circuit schematic including qubit, resonator, and the p-SQUID. Black symbols denote the model parameters. b Potential energy of an RF-SQUID as a function of the flux in its loop $\Phi_{\mathrm{pSQ}}$. Transitions between wells correspond to a change of $\Phi_{\mathrm{pSQ}}$ by $\approx 1\Phi_{0}$. c RF-SQUID flux $\Phi_{\mathrm{pSQ}}$ vs. the applied external flux $\Phi_{\mathrm{ext}}$ for the different sweep directions (arrows). d RF-SQUID resonance frequency $f_{\mathrm{pSQ}}$ vs. $\Phi_{\mathrm{ext}}$. Dashed arrows indicate inter-well transitions. e The qubit resonance frequency $f_{\mathrm{Q}}$ (bright pixels), measured vs. the applied magnetic field $B_{\parallel}$, shows a sawtooth behaviour due to the discretized flux induced from the p-SQUID.
Figure 2a shows the expected periodic oscillations in the resonator’s dispersive frequency shift when the qubit is tuned between its lowest and highest (at 4.125 GHz) resonance frequencies according to Eqs. (1) and (2) by sweeping the magnetic field $B_{\parallel}$. We note that $B_{\parallel}$ should nominally be oriented perpendicular to the plane of the qubit’s SQUID loop and therefore not couple to $\Phi_{\mathrm{Q}}$. Accordingly, we use these data to characterize the misalignment of the solenoid (see App. D for further details).
@@ Results
The model agrees excellently with both experiments for the same set of fit parameters (red lines in Figs. 4a and b). We note that the data obtained near maxima of the qubit frequency is distorted due to the small detuning of qubit and resonator, and has therefore been omitted from the fits. Since the least-squares fit involved 8 free parameters (see App. A), large uncertainties in the fitting values don’t allow us to obtain detailed information on the origin of the parasitic junction that gave rise to the p-SQUID. However, the observed hysteresis clearly reveals the p-SQUID’s multi-well potential associated with [9] $\beta_{L,\mathrm{pSQ}}=2\pi\,L_{\mathrm{\mathrm{pSQ}}}\cdot I_{c,\mathrm{pSQ}}/\Phi_{0}>1$. This places the p-SQUID in the parameter regime of flux-biased phase qubits [22, 23, 24]. Estimating the p-SQUID’s inductance to $L_{\mathrm{pSQ}}\approx 0.5\,$nH (see App. C), this corresponds to a critical current of the parasitic junction of $I_{\mathrm{c,pSQ}}\gtrsim 1\mu$A. Aiming to identify the flawed bond after the sample was warmed up, the bond wire bridges were touched and heavily bent using the micro-manipulator of a bonding station. However, all bonds appeared stable and did not disconnect easily. We note that it has been proposed to utilize the discreteness of RF-SQUID flux states to provide a stable flux-bias source for qubits [25, 26].
Discussion
@@
While it has been known that stray Josephson junctions between thin-film qubit electrodes are a source of qubit decoherence [27, 17], our finding sheds light on related issues in the remaining circuitry and packaging. In our experiment, two out of four qubits on the chip were inoperable due to their strong interference with the parasitic RF-SQUID formed by a faulty chip-to-chip wirebond. Related issues may arise for instance from detrimental modes in flux bias lines that were terminated by a junction in a flawed wirebond. Problems with parasitic junctions due to flawed airbridges, bond wires, or bump-pads in flip-chip architectures[28], may rise in importance as quantum processors increase in size and complexity. The observed strong DC-magnetic and AC-electric interactions of parasitic RFSQUIDs can furthermore spoil the functionality of various other superconducting devices [29] such as for instance parametric amplifiers[30, 31], RSFQ-logic [32], MKID photon detectors[33], and passive resonator-based microwave filters [34].
@@ Acknowledgements.
We acknowledge the contributions of Alexander Bilmes to the experimental setup and qubit fabrication. We thank Hannes Rotzinger for fruitful discussions and assistance in the laboratory. BB acknowledges help of Konstantin Händel with sample fabrication. This work was funded by Google, which we gratefully acknowledge.
@@ Appendix A Fit to the model
$g_{\mathrm{pSQ},\mathrm{r}}$ 10.2 MHz coupling strength pSQUID-resonator
$g_{\mathrm{pSQ},\mathrm{q}}$ 340 MHz coupling strength pSQUID-qubit
$\phi_{slope}^{Q}$ 0.19 $\Phi_{0}$ $\Phi_{\mathrm{Q}}$induced between well jumps
$\phi_{qjump}^{Q}$ 0.12 $\Phi_{0}$ $\Phi_{\mathrm{Q}}$jump at well jump
$L_{\mathrm{pSQ}}$ 1.23 nH p-SQUID inductance
$C_{\mathrm{pSQ}}$ 3.1 pF p-SQUID capacitance
$I_{c,\mathrm{pSQ}}$ 1.53 $\mu$A p-SQUID critical current
@@
Figure 5: Resonator resonance frequency $f_{\mathrm{r}}$ vs. $B_{\parallel}$ obtained from the Hamiltonian Eq. 3 when different coupling parameters are set to zero. Without a direct transversal coupling between p-SQUID and the resonator (yellow line), the sharp dips near the minimum of $f_{\mathrm{r}}$ are absent. The blue line shows only the dips induced by this direct interaction when both the magnetic coupling $M_{pSQ,q}$ and AC coupling $g_{SQ,q}$ between qubit and p-SQUID  are set to zero.
Appendix C Bond wire inductance
Superconducting bond wires are typically made from aluminum wires of 20 to 25 $\mu$m diameter. A common rule-of-thumb is that 1mm of bond wire has an inductance of 1 nH. We compare this rule with the textbook formula [35]
$L=\frac{\mu_{0}}{2\pi}l\left(\log\left(\frac{2l}{r}\right)-1\right)$ (4)
where $l$ is the wire length and $r$ the wire radius as shown in Figure 6 b. Measuring the lengths of the bond wires on our sample using optical microscopy from different angles (see Fig. 6a), we estimate their inductances to range from 240pH to 380 pH.
@@ Appendix D Field alignment and p-SQUID geometry
The effective p-SQUID area is rather small due to gradiometer effects in the four parallel wirebond connections, see Fig. 1 and Fig. 7.

### v2

@@ II Experiment
The qubit transition frequency $f_{\mathrm{Q}}\$depends on the flux $\Phi_{\mathrm{Q}}$ threading its DC-SQUID and can be approximated by [12]
@@
Figure 3: a Circuit schematic including qubit, resonator, and the p-SQUID. Black symbols denote the model parameters. b Potential energy of an RF-SQUID as a function of the flux in its loop $\Phi_{\mathrm{pSQ}}$. Transitions between wells correspond to a change of $\Phi_{\mathrm{pSQ}}$ by $\approx 1\Phi_{0}$. c RF-SQUID flux $\Phi_{\mathrm{pSQ}}$ vs. the applied external flux $\Phi_{\mathrm{ext}}$ for the different sweep directions (arrows). d RF-SQUID resonance frequency $f_{\mathrm{pSQ}}$ vs. $\Phi_{\mathrm{ext}}$. Blue and red lines indicate different sweep directions. e The qubit resonance frequency $f_{\mathrm{Q}}$ (bright pixels), measured vs. the applied magnetic field $B_{\parallel}$, shows a sawtooth behaviour due to the discretized flux induced from the p-SQUID.
Figure 2a shows the expected periodic oscillations in the resonator’s dispersive frequency shift when the qubit is tuned between its lowest and highest (at 4.5662 GHz $\pm$ 200 kHz) resonance frequencies according to Eqs. (1) and (2) by sweeping the magnetic field $B_{\parallel}$. We note that $B_{\parallel}$ should nominally be oriented perpendicular to the plane of the qubit’s SQUID loop and therefore not couple to $\Phi_{\mathrm{Q}}$. Accordingly, we use these data to characterize the misalignment of the solenoid (see App. D for further details).
@@ III Model
The model agrees excellently with both experiments for the same set of fit parameters (red lines in Figs. 4a and b). We note that the data obtained near maxima of the qubit frequency is distorted due to the small detuning of qubit and resonator, and has therefore been omitted from the fits. Since the least-squares fit involved 8 free parameters (see App. A), large uncertainties in the fitting values don’t allow us to obtain detailed information on the origin of the parasitic junction that gave rise to the p-SQUID. However, the observed hysteresis clearly reveals the p-SQUID’s multi-well potential associated with [9] $\beta_{L,\mathrm{pSQ}}=2\pi\,L_{\mathrm{\mathrm{pSQ}}}\cdot I_{c,\mathrm{pSQ}}/\Phi_{0}>1$. This places the p-SQUID in the parameter regime of flux-biased phase qubits [22, 23, 24]. We estimate the possible ranges of p-SQUID parameters by fixing $\beta_{L,\mathrm{pSQ}}$ to a value of 11.3 which is compatible with the observed hysteresis. The restricted fit (see Appendix C) achieves good agreement with the data when $L_{\mathrm{pSQ}}$ ranges between $0.1$ and 1.7 nH, corresponding to critical currents of $I_{c,\mathrm{pSQ}}$ between $80\mu A$ and $20\mu A$ and capacitances $C_{\mathrm{pSQ}}$ of 30pF to 5 pF. While the best-fitting inductance of $L_{\mathrm{pSQ}}\approx 474$ pH is well in accordance with the wirebonds’ inductances, the capacitance and critical current of the parasitic junction can not be predicted without exact knowledge of its geometry since it depends strongly on the dielectric material (such as photoresist residual) and its thickness.
Aiming to identify the flawed bond after the sample was warmed up, the bond wire bridges were touched and heavily bent using the micro-manipulator of a bonding station. However, all bonds appeared stable and did not disconnect easily. We note that it has been proposed to utilize the discreteness of RF-SQUID flux states to provide a stable flux-bias source for qubits [25, 26].
IV Discussion
The observed anomaly of resonator and qubit resonance frequencies in applied magnetic fields is fully explained by their interactions with a parasitic RF-SQUID. The notion that a faulty wirebond contains a parasitic tunnel junction is supported by in- and out-of-plane magnetic field sweeps which confirm the three-dimensional form of the p-SQUID’s loop. A further confirmation is the fact that the anomaly disappeared once the wirebonds were removed. Other possible origins of flux irregularities, such as the motion of trapped vortices [27], would not result in the observed strictly periodic signatures and observed dispersive resonance frequency shifts.
@@ IV Discussion
While it has been known that stray Josephson junctions between thin-film qubit electrodes are a source of qubit decoherence [28, 17], our finding sheds light on related issues in the remaining circuitry and packaging. In our experiment, two out of four qubits on the chip were inoperable due to their strong interference with the parasitic RF-SQUID formed by a faulty chip-to-chip wirebond.
The signature of a parasitic SQUID was so far only observed in a single experiment. To determine how common this problem actually is will require a systematic study on many samples, in which a large range of applied magnetic fields is explored to search for anomalies such as avoided level crossings in resonator and qubit spectroscopy. It can be seen as a coincidence that the parameters of the p-SQUID placed it into a regime where its interactions with qubit and resonator could be clearly observed. In more typical cases, the p-SQUID could be further detuned and induce only slight frequency shifts in a circuit that may remain unnoticed. Similarly, p-SQUIDs may be overlooked when only on-chip qubit flux bias lines are used that only slightly tune the p-SQUID’s resonance frequency, while their parasitic modes would still contribute to decoherence and flux noise.
Related issues may arise for instance from detrimental modes in flux bias lines that were terminated by a junction in a flawed wirebond. Problems with parasitic junctions due to flawed airbridges, bond wires, or bump-pads in flip-chip architectures[29], may rise in importance as quantum processors increase in size and complexity. The observed strong DC-magnetic and AC-electric interactions of parasitic RFSQUIDs can furthermore spoil the functionality of various other superconducting devices [30] such as for instance parametric amplifiers[31, 32], RSFQ-logic [33], MKID photon detectors[34], and passive resonator-based microwave filters [35].
@@ Acknowledgements.
We acknowledge the contributions of Alexander Bilmes to the experimental setup and qubit fabrication. We thank Hannes Rotzinger for fruitful discussions and vital help in the laboratory and clean-room. BB acknowledges help of Konstantin Händel with sample fabrication. This work was funded by Google, which we gratefully acknowledge.
@@ Appendix A Fit to the model
$g_{\mathrm{pSQ},\mathrm{r}}$ 9.9 MHz coupling strength pSQUID-resonator
$g_{\mathrm{pSQ},\mathrm{q}}$ 70.3 MHz coupling strength pSQUID-qubit
$\phi_{slope}^{Q}$ 0.177 $\Phi_{0}$ $\Phi_{\mathrm{Q}}$induced between well jumps
$\phi_{qjump}^{Q}$ 0.116 $\Phi_{0}$ $\Phi_{\mathrm{Q}}$jump at well jump
$L_{\mathrm{pSQ}}$ 474 pH p-SQUID inductance
$C_{\mathrm{pSQ}}$ 10.3 pF p-SQUID capacitance
[…]
@@
To estimate the range of pSQUID parameters which result in reasonable agreement with the data, we fixed the parameter $\beta_{L}=2\pi L_{\mathrm{pSQ}}I_{\mathrm{C,pSQ}}/\Phi_{0}$ to the value of 11.3 to ensure that the p-SQUID hysteresis matches the data shown in Fig. 2. A value of the pSQUID inductance $L_{\mathrm{pSQ}}$ was then chosen, and the pJJ’s critical current $I_{\mathrm{C,pSQ}}$ was calculated accordingly. The other parameters listed in Table 1 were kept as free fitting parameters.
Figure 6b shows the deviation from the data in the form of the root mean square error of the fit. The fit quality remains rather constant for a wide range of $L_{\mathrm{pSQ}}$ between 0.1-1.7 nH. For higher values, the fit quickly becomes worse as it becomes impossible to fulfill the three boundary conditions imposed by the p-SQUID-hysteresis, the fixed $L_{\mathrm{pSQ}}$, and the requirement that the pSQUID’s resonance frequency must always exceed that of the readout resonator.
[…]
@@
Figure 5: Resonator resonance frequency $f_{\mathrm{r}}$ vs. $B_{\parallel}$ obtained from the Hamiltonian Eq. 3 when different coupling parameters are set to zero. Without a direct transversal coupling between p-SQUID and the resonator (yellow line), the sharp dips near the minimum of $f_{\mathrm{r}}$ are absent. The blue line shows the dips induced by this direct interaction.
Appendix C Estimating the pSQUID parameters
To estimate the range of realistic values for the inductance of the pSQUID, we consider the geometry of the wirebonds shown in Fig. 1. From optical microscopy as shown in Fig. 8a, the lengths of the four wirebonds is measured as 381µm, 466µm, 466µm, and 546µm, respectively. Using the textbook formula [36]
$L=\frac{\mu_{0}}{2\pi}l\left(\log\left(\frac{2l}{r}\right)-1\right),$ (4)
where $r$ is the wire radius and $l$ is its length, we estimate inductances of 237pH to 379pH for our aluminum bond wires which have a diameter of 25µm. In Fig. 8b, the inductance per length is plotted also for other wire diameters, and compared to a common rule of thumb which estimates that a 1mm-long bond wire has a typical inductance of 1 nH.
To obtain a lower limit for the range of reasonable pSQUID inductance, we assume that the parasitic junction is contained in the shortest bond wire, and that the pSQUID loop is closed by the parallel connection of the remaining three wirebonds. The upper limit is estimated by assuming that only the two longest bond wires are involved in the loop formation. This results in an estimated range of $347\mathrm{pH}\lesssim L_{\mathrm{pSQ}}\lesssim 688$ pH, which is well in accordance with the best fitting parameter $L_{\mathrm{pSQ}}$ of 474 pH. The limits are indicated by vertical dashed lines in Fig. 6.
Figure 6: Top panel: Fitting values of $C_{\mathrm{pSQ}}$ and $I_{c,\mathrm{pSQ}}$ as a function of $L_{\mathrm{pSQ}}$ for fixed $\beta_{L}\equiv 2\pi LI_{C}/\Phi_{0}=11.3$ (as determined by a measurement of the p-SQUID’s hysteresis). The bottom panel shows the root mean squared error of the fit, which remains similarly small in a wide range of $L_{\mathrm{pSQ}}$. The vertical dashed lines in the top panel indicate the realistic range of bond wire inductances as described in the text.
For the pSQUID capacitance, an upper limit can be estimated from the average area of a wirebond foot that is measured to about $A_{\mathrm{foot}}\approx(62\,\mu m)^{2}=3840\,\mu\mathrm{m}^{2}$. Taking d=2d=2\,nm as a minimal thickness of the capacitor dielectric (so that the whole bond wire foot would form a tunnel junction) and a relative permittivity $\varepsilon_{r}\approx 10$ for AlOx, we obtain $C_{\mathrm{pSQ}}=\epsilon_{0}\epsilon_{r}A_{\mathrm{foot}}/d=$170 pF which is much larger than the fitted $C_{\mathrm{pSQ}}\approx 10$pF. Realistically, the whole bond foot area must contribute to $C_{\mathrm{pSQ}}$. The discrepancy between this large area and relatively small (fitting) capacitance can be explained by a larger thickness and lower permittivity of the dielectric under the faulty bonding wire foot. For example, assuming that the bond foot area is contaminated by a 10 nm-thick layer of residual PMMA photoresist ($\varepsilon_{r}\approx 3$) results in a capacitance of 10 pF that corresponds to the best fitting value.
To reconcile the large bond foot area with the small capacitance (requiring a thick dielectric) and formation of a tunnel junction (requiring a tunnel barrier thickness of $\lesssim 3$nm), one can assume that the thickness of the dielectric under the wirebond foot is non-uniform, possibly due to photoresist residuals. Figure 7 illustrates this situation in a cross-section through the wirebond foot. The actual contamination of our sample surface is unknown. In ultrasonic wirebonding, the amount of intermetallic formed at the bonded interface depends strongly on the bonding parameters such as ultrasound frequency, amplitude, and duration [37], so that inadequate settings could have played a role in the junction formation.
Figure 7: Sketch of the cross-section through the foot of a wirebond (not to scale), illustrating thickness variations of the native AlOx on the ground plane and a layer of contamination such as photoresist residual. The active junction area where Cooper pair tunneling occurs may be limited to small regions of least dielectric thickness in addition to pinholes.
[…]
@@ Appendix D Field alignment and p-SQUID geometry
The effective p-SQUID area is rather small due to gradiometer effects in the four parallel wirebond connections as shown in Fig. 1 and the illustration in Fig. 9.
@@
Appendix E Sample fabrication
[…]
This aluminum ground plane is then structured using optical lithography. For this purpose, S1805 photoresist is applied and patterned using a mask-aligner and AZ Developer, followed by Al etching in an inductively coupled plasma. Afterwards, the resist is removed and the wafer cleaned with DMSO, 2-propanol and an oxygen plasma.
The Josephson junctions and leads were fabricated using a modified version of the in-situ bandaged Niemeyer-Dolan technique that avoids the formation of stray junctions [38]. For this, a bi-layer of $250\text{\,}\mathrm{nm}$ A4 (PMMA) on top of $900\text{\,}\mathrm{nm}$ EL-13 (MMA) was applied to the sample and patterned using an electron beam writing system and developed with a mixture of 2-propanol and bidistilled water. Resist residuals in the trenches were removed using an oxygen plasma. Afterwards, the Josephson junctions and leads were deposited in the PLASYS MEB 550 S at a base pressure of $5\times 10^{-8}$\mathrm{mbar}$\$.
Deviating from conventional fabrication, the here-used samples employed thicker-than-usual tunnel barriers. These were realized by first oxidizing the deposited junction bottom electrodes at a static pressure of $130\text{\,}\mathrm{mbar}$ for $20\text{\,}\mathrm{min}$. To further thicken the barrier, an additional aluminium layer of about $0.1\text{\,}\mathrm{nm}$ thickness was deposited and thoroughly oxidized, followed by deposition of the junction’s top electrode. After protective oxidation at $30\text{\,}\mathrm{mbar}$ for $10\text{\,}\mathrm{min}$, a lift-off was performed using DMSO and the sample is cleaned in 2-propanol and an ultrasonic bath. The resulting tunnel junctions had critical current densities of about 2 A/cm2. A thorough discussion of junction fabrication and qubit coherence will be the focus of an upcoming work [39].
Appendix F Experimental setup
The sample was measured in a dry dilution refrigerator at a base temperature of about 25 mK. The readout resonator frequency was measured with a vector network analyzer via a transmission line that had total attenuation of 80 dB distributed over several temperature stages, and was equipped with band-pass and infrared filters. After interaction with the resonator, the signal was isolated by means of two circulators and additional band-pass and infrared filters from a HEMT amplifier at the 4K stage, and further amplified by 66 dB at room temperature.